hydrogen gas effuses through a porous container 8.97-times faster than an unknown gas. Estimate the molar mass of the unknown gas. 


1
Expert's answer
2021-05-20T07:53:59-0400

MM(H2) = 2.016 g/mol

"\frac{Rate_A}{Rate_B} = \sqrt{\frac{MM_B}{MM_A}} \\

8.97 = \sqrt{ \frac{MM_B}{2} } \\

(8.97)^2 = \frac{MM_B}{2.016} \\

MM_B = (8.97)^2 \times 2.016 = 162.21 \;g/mol"

Answer: 162.21 g/mol